OpenStudy (anonymous):

Simplify the expression. 34 + 12 ÷ 3 - (1-9) a.)75 b.)39 c.)93 d.)21

OpenStudy (anonymous):

the 4 is a exponent .

OpenStudy (anonymous):

34+4-(-8) 34+4+8=46 Are you sure is correct answer?

OpenStudy (anonymous):

Nancy, the 4's an exponent...

OpenStudy (anonymous):

hibarbie, use the '^' symbol for an exponent.

OpenStudy (anonymous):

Here, you'd write 3^4

OpenStudy (anonymous):

Okay?

OpenStudy (anonymous):

\[3^4+12\div3-(1-9)\] Remember the order of operations: Brackets, Exponents, Division, Multiplication, Addition, Subtraction. When there is a conflict of what to do first, do them in order as they appear from left to right. So with this, that would be brackets first (1-9 equals -8), then exponents (3^4)=81, then division (12/3)=4. So what you're left with is 81+4+8 (remember that subtracting a negative number is the same as adding the positive) which would be 93.